### v1

@@ 2.2 DiffSpectra generates valid and stable molecular structures
As summarized in Table 1, our proposed DiffSpectra consistently achieves strong results in generating chemically valid and stable molecular structures under spectrum-conditioned settings. In the 2D evaluation, DiffSpectra attains 99.9% atom stability and 98.2% molecular stability, demonstrating performance comparable to the training data distribution, while outperforming or matching general molecular generative models such as CDGS [36] and JODO [37]. Moreover, DiffSpectra maintains high structural uniqueness and novelty, with V&U (valid and unique) at 96.8% and V&U&N (valid, unique, and novel) at 93.7%, significantly surpassing prior methods. In distribution-based metrics, DiffSpectra achieves the lowest FCD, indicating a close distributional match to the reference test set. It also reaches the best SNN, Frag, and Scaf scores, demonstrating a strong capacity to cover diverse chemical substructures and scaffolds while maintaining validity.
In the 3D evaluation, DiffSpectra demonstrates competitive geometric stability, with atom stability and molecular stability again comparable to the training set. Notably, DiffSpectra achieves the best angle and dihedral MMDs, suggesting accurate recovery of bond angles and torsional angles, and achieves a bond distance MMD comparable to the top-performing baselines. This highlights DiffSpectra ’s effectiveness in reconstructing not only 2D connectivity but also realistic 3D molecular conformations.
The Train row in Table 1 reports the statistics of the training dataset as a reference distribution. For some metrics that measure similarity to the test set, such as SNN, Frag, Scaf, or geometric MMDs, the training data does not necessarily achieve the best possible scores, since the test distribution may differ from the training distribution. All other baselines listed in Table 1, including CDGS [36], JODO [37], EDM [41], and others, are unconditional molecular generative models, which sample from the general molecular space without explicit external guidance. In contrast, DiffSpectra is specifically designed for molecular structure elucidation conditioned on spectra. The use of spectral signals as conditions provides external structural priors, enabling the model to generate results that better match the test set distribution. Therefore, the higher V&U, V&U&N, SNN, Frag, Scaf, and MMD metrics achieved by DiffSpectra compared to unconditional models are expected, as the conditioning naturally constrains the generation process toward structures consistent with the observed test samples.
Figure 2: Visualization of molecular structure elucidation results using DiffSpectra under different configurations. We compare single-spectrum inputs (IR, Raman, UV-Vis), multi-modal spectra, and the effect of incorporating the pre-trained SpecFormer. Ground-truth structures are shown on the left for reference.
Table 2: Structure elucidation performance of DiffSpectra on the QM9S dataset. We compare two configurations: one using a pre-trained SpecFormer as the spectral condition encoder, and one using an untrained SpecFormer. Reported metrics include exact structure recovery, graph overlap, fingerprint-based similarities, fragment-level similarity, and functional group similarity.
Model Pre-trained SpecFormer ACC@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$uparrow $\operatorname{CosSim}_{\mathrm{MG}}$uparrow $\operatorname{TaniSim}_{\mathrm{MA}}$uparrow FraggleSim $\uparrow$ FGSim $\uparrow$ DiffSpectra 16.01% 1.3552 0.7837 0.8421 0.9227 0.9481 0.9618 14.11% 1.7795 0.7205 0.7938 0.8924 0.9383 0.9490
@@ 2.3 DiffSpectra accurately elucidates molecular structures from spectra
As shown in Table 2 and Figure 2, DiffSpectra demonstrates strong performance in the challenging task of molecular structure elucidation from spectra. First, the top-1 accuracy (ACC@1) reaches 16.01%, meaning that the model is able to recover the exact target structure among its predictions in a nontrivial fraction of cases. Beyond strict exact matches, the average MCES score is 1.3552, indicating that even when the model fails to generate an exact structure, it still preserves a substantial portion of the molecular graph connectivity.
@@
In terms of fingerprint-based similarity metrics, DiffSpectra achieves a Tanimoto similarity of 0.7837 and a cosine similarity of 0.8421 over Morgan fingerprints, reflecting high agreement on local structural features. The Tanimoto similarity over MACCS keys is even higher at 0.9227, highlighting good coverage of predefined chemical substructures. These results suggest that DiffSpectra captures the key functional groups and local motifs characteristic of the target molecules.
@@
Furthermore, the fragment-based Fraggle similarity reaches 0.9481, showing that large chemically meaningful fragments are well recovered. The functional group similarity (FGSim) is also high at 0.9618, confirming that the predicted molecules retain nearly all functional group types present in the ground-truth structures. Overall, these metrics collectively indicate that DiffSpectra is able to accurately elucidate molecular structures from spectra, recovering both global connectivity and local functional features with high fidelity.
@@ 2.4 Pre-trained SpecFormer facilitates more accurate structure elucidation
To investigate the effect of pre-training the spectrum encoder, we conduct an ablation study comparing DiffSpectra with and without a pre-trained SpecFormer module, as summarized in Table 2. When equipped with the pre-trained SpecFormer, DiffSpectra achieves a top-1 accuracy of 16.01%, compared to 14.11% without pre-training, indicating a clear improvement in correctly recovering the ground-truth molecular structures from spectra. Similarly, the MCES decreases from 1.7795 to 1.3552, suggesting that the molecular graphs predicted with pre-trained SpecFormer retain more of the correct connectivity of the target structures.
In addition, the fingerprint-based similarity scores also improve with pre-training. For example, the Tanimoto similarity over Morgan fingerprints rises from 0.7205 to 0.7837, and cosine similarity improves from 0.7938 to 0.8421. The Tanimoto similarity over MACCS keys also increases from 0.8924 to 0.9227. These gains indicate that the molecular structures generated with pre-trained SpecFormer more accurately capture both local substructural patterns and functional group motifs. Moreover, the Fraggle similarity improves from 0.9383 to 0.9481, and FGSim from 0.9490 to 0.9618, confirming that larger chemical fragments and functional groups are better preserved.
@@
Table 3: Effect of spectral modalities on structure elucidation performance. We report DiffSpectra’s results on the QM9S dataset using different types of spectra as conditional input, including IR, Raman, UV-Vis, and their combination. Metrics include exact structure recovery, graph overlap, fingerprint-based similarities, fragment-level similarity, and functional group similarity.
Model Spectral Modalities ACC@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$uparrow $\operatorname{CosSim}_{\mathrm{MG}}$uparrow $\operatorname{TaniSim}_{\mathrm{MA}}$uparrow FraggleSim $\uparrow$ FGSim $\uparrow$ DiffSpectra All Spectra 14.11% 1.7795 0.7205 0.7938 0.8924 0.9383 0.9490 Only IR 10.97% 2.4812 0.6246 0.7188 0.8460 0.9197 0.9269 Only Raman 12.51% 2.1708 0.6778 0.7612 0.8701 0.9343 0.9315 Only UV-Vis 0.10% 8.7909 0.1556 0.2625 0.3634 0.5581 0.4567
@@ 2.5 Multi-modal spectra outperform single-modality spectra for structure elucidation
Among individual modalities, Raman spectra alone outperform IR and UV-Vis, with ACC@1 of 12.51% and reasonable similarity scores. IR spectra alone still provide useful chemical information, achieving 10.97% top-1 accuracy, while UV-Vis alone performs poorly (0.1% top-1 accuracy), reflecting its limited ability to uniquely identify molecular structures in the QM9S dataset. Overall, these results highlight that combining multiple spectra modalities provides complementary structural priors, enabling DiffSpectra to generate molecular structures with higher fidelity and consistency to ground-truth targets.
@@ 2.6 Sampling multiple candidates improves structural hit accuracy
Figure 3: Accuracy@$K$ with increasing number of sampled candidates. We evaluate the top-$K$ accuracy (ACC@$K$) as the number of generated candidates $K$ increases, comparing different variants of our model. Across all settings, Accuracy@$K$ consistently improves with larger $K$, confirming that multiple sampling significantly increases the chance of recovering the correct molecular structure.
@@
Due to the inherently stochastic nature of diffusion models, each sampling process can produce a distinct yet plausible molecular structure. While a single sample may not always match the ground-truth structure exactly, generating multiple candidates increases the likelihood that at least one of them will align with the correct structure. This motivates the use of Top-$K$ Accuracy (ACC@$K$) as a more comprehensive evaluation metric, which is formally defined in Section C.2.
@@
We present in Figure 3 the performance of different model variants under varying numbers of samples. Specifically, ACC@$K$ measures the proportion of test molecules for which the correct structure appears among the top $K$ generated candidates. For instance, the ACC@1 reflects the exact match rate under single-sample generation, while ACC@$K$ ($K>1$) demonstrate that allowing multiple guesses significantly boosts hit probability. Across all model variants, we observe a consistent upward trend in Accuracy@$K$ as $K$ increases. As $K$ increases, the model has more opportunities to generate a correct structure among its top-$K$ outputs. For example, the full model with all spectra and pre-trained SpecFormer improves from 16.01% at $K=1$ to 96.86% at $K=20$, showing that even a small sampling budget can significantly improve overall accuracy.
These results highlight an important property of diffusion-based molecular elucidation: even when the model may not always generate the correct structure in one shot, it maintains a strong ability to place the correct structure within a small number of plausible candidates. This aligns well with practical use cases, where a ranked list of likely structures can be provided for downstream validation or experimental testing. The effectiveness of multi-sample inference further demonstrates how spectral conditioning constrains the generative space, enabling the model to consistently place the correct structure among top candidates.
@@
Table 4: Evaluation of different SE(3) equivariance strategies for DiffSpectra. We compare a model-based equivariant architecture to data-based equivariance approaches, with or without data augmentation, on molecular structure elucidation metrics. Here, “data aug” refers to data augmentation.
@@
Model SE(3) Equivariance ACC@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$uparrow $\operatorname{CosSim}_{\mathrm{MG}}$uparrow $\operatorname{TaniSim}_{\mathrm{MA}}$uparrow FraggleSim $\uparrow$ FGSim $\uparrow$ DiffSpectra model-based 14.11% 1.7795 0.7205 0.7938 0.8924 0.9383 0.9490 data-based (w/ data aug) 12.98% 1.8785 0.8882 0.7877 0.8882 0.9389 0.9456 data-based (w/o data aug) 7.47% 3.6036 0.5117 0.6294 0.7575 0.8665 0.8607
@@ 2.7 Comparing model-based and data-based SE(3) equivariance strategies
Since our model performs diffusion generation on molecular 3D structures, it is essential to ensure SE(3) equivariance so that the model’s predictions are consistent under rigid-body transformations of the input. To this end, we design and evaluate two strategies for achieving SE(3) equivariance: model-based equivariance and data-based equivariance. In the model-based design, geometric inductive biases are explicitly incorporated into the network architecture, for example through pairwise distance encoding and equivariant coordinate update mechanisms, ensuring equivariant behavior. In contrast, the data-based design does not include SE(3)-equivariant inductive biases in the model architecture. Instead, it relies on random rotation and translation data augmentation during training, encouraging the model to empirically learn equivariance from data.
@@
As shown in Table 4, the model-based SE(3)-equivariant implementation of DiffSpectra achieves superior performance across most metrics. It attains a top-1 accuracy of 14.11% and the lowest MCES (1.7795), indicating stronger recovery of molecular graph connectivity. The data-based approach with data augmentation achieves slightly lower top-1 accuracy (12.98%) but shows competitive fingerprint-based and functional group similarities, with TanimotoMG, FraggleSim, and FGSim comparable to the model-based design. Notably, when data augmentation is removed from the data-based design, performance degrades substantially across all metrics, confirming the critical role of data augmentation in enabling the model to generalize over molecular rotations and translations. Overall, these results demonstrate that while data-based equivariance with augmentation can help achieve some level of SE(3)-aware behavior, explicitly enforcing geometric inductive biases through model-based equivariance is a more robust and effective strategy for molecular structure elucidation from spectra.

### v2

@@ 2.2 DiffSpectra generates valid and stable molecular structures
As summarized in Table 1, our proposed DiffSpectra consistently achieves strong performance in generating chemically valid and stable molecular structures under spectra-conditioned settings. It attains 99.9% atom stability and 98.2% molecular stability, demonstrating performance comparable to that of the training data distribution, and outperforming or matching general unconditional molecular generation models such as CDGS [44] and JODO [45]. DiffSpectra also maintains high structural uniqueness and novelty, with V&U (valid and unique) at 96.8% and V&U&N (valid, unique, and novel) at 93.7%, significantly surpassing prior methods. On distribution-based metrics, DiffSpectra achieves the lowest Fréchet ChemNet Distance (FCD), indicating a close alignment with the test distribution, and the best SNN, Frag, and Scaf scores, demonstrating strong coverage of diverse chemical substructures and scaffolds while maintaining validity.
The Training Set row in Table 1 reports the statistics of the training set as the reference distribution. For metrics that measure similarity to the test set, such as SNN, Frag, or Scaf, the training data do not necessarily achieve the highest scores, since the test distribution may differ from the training distribution. All other baselines listed in Table 1 are unconditional molecular generation models that sample the general molecular space without explicit conditions. In contrast, DiffSpectra is specifically designed for molecular-structure elucidation conditioned on spectra. Spectral signals provide external structural priors, enabling the model to generate structures that better match the test set distribution. Therefore, the higher V&U, V&U&N, FCD, SNN, Frag, and Scaf metrics achieved by DiffSpectra compared to unconditional models are expected, as the conditioning naturally constrains the generation process toward structures consistent with the observed test samples.
@@ 2.3 DiffSpectra accurately elucidates molecular structures from spectra
As shown in Figure 2(A), Table 2, and Figure 3, DiffSpectra demonstrates strong performance on the challenging task of molecular structure elucidation from spectra. The top-1 accuracy (Acc@1) reaches 40.76%, indicating that the model can recover the exact target structure among its predictions in a non-trivial fraction of cases. Beyond strict exact matches, the average MCES score is 1.3337, suggesting that even when the model fails to generate an exact structure, it still preserves a substantial portion of the molecular graph connectivity.
Figure 2: Quantitative comparison of DiffSpectra and its variants on molecular structure elucidation tasks. a, Structure elucidation performance of DiffSpectra under different configurations. We compared the performance when using a pre-trained SpecFormer as the spectral condition encoder versus an un-pretrained SpecFormer, as well as when using multi-modal spectra versus individual spectra (IR, Raman, UV-Vis). Reported metrics include exact top-1 accuracy, MCES, and several similarity-based metrics. b, Evaluation of elucidated 3D molecular structures. For each generated 3D structure, we performed atom-to-atom mapping against the corresponding ground-truth 3D structure and further computed the RMSD. When using a pre-trained SpecFormer with multi-modal spectra as input, DiffSpectra achieves the best elucidation performance, showing the highest atom mapping accuracy and the lowest RMSD. c, Accuracy@$K$ with an increasing number of sampled candidates. We report top-$K$ accuracy as the number of generated candidates $K$ increases. Across all settings, Accuracy@$K$ consistently improves with larger $K$, confirming that multiple sampling significantly increases the likelihood of recovering the correct molecular structure. d, Distribution of similarity-based metrics on the test set, illustrated using violin and box plots. The red line in each box plot denotes the median. e and f, Occurrence of functional groups in the test molecules and the corresponding elucidation performance of DiffSpectra. To account for the imbalance in functional group occurrence, we computed the ROC-AUC score for each group based on whether DiffSpectra correctly identified its presence or absence in each molecule.
@@
In terms of fingerprint-based similarity metrics, DiffSpectra achieves a Tanimoto similarity of 0.7875 and a cosine similarity of 0.8450 over Morgan fingerprints, reflecting high agreement with local structural features. The Tanimoto similarity over MACCS keys is even higher at 0.9262, highlighting broad coverage of predefined chemical substructures. These results suggest that DiffSpectra effectively captures the key functional groups and local motifs characteristic of the target molecules. Furthermore, the fragment-based Fraggle similarity reaches 0.9504, showing that large chemically meaningful fragments are well recovered. The functional group similarity (FGSim) is also high at 0.9652, confirming that the predicted molecules retain nearly all functional group types present in the ground-truth structures.
To further verify the quality of the elucidated 3D molecular structures, we performed atom mapping against the corresponding ground-truth 3D structures using the Hungarian algorithm and computed the root mean square deviation (RMSD) in Å. As shown in Figure 2(b), when using a pre-trained SpecFormer with multi-modal spectra as input, DiffSpectra achieves the best elucidation performance, reaching 98.25% atom mapping accuracy and an RMSD of 1.651. These results highlight the ability of our model to elucidate high-quality 3D structures with accurate atom types and precise coordinates.
We examined the detailed distributions of similarity-based metrics on the test set, as shown in Figure 2(d). These metrics provide quantitative measures of how closely the elucidated structures resemble their corresponding ground-truth molecules. Across all metrics, DiffSpectra, when equipped with a pre-trained SpecFormer and utilizing all spectral modalities as input, exhibits the most favorable distributions. This trend aligns well with the overall performance trends reported in Figure 2(a).
We further analyzed the elucidation performance with respect to functional groups, as illustrated in Figure 2(e–f). Figure 2(e) presents the distribution of functional groups in the test molecules, showing a pronounced imbalance in their occurrence frequency. To address this imbalance, we evaluated model performance using the ROC-AUC score for each group, which reflects whether DiffSpectra correctly predicts the presence or absence of a given functional group in each molecule. As shown in Figure 2(f), DiffSpectra achieves consistently high ROC-AUC values across diverse groups, demonstrating its ability to maintain reliable elucidation performance even under imbalanced data distributions.
@@
Overall, these metrics collectively indicate that DiffSpectra can accurately elucidate molecular structures from spectra, recovering not only global connectivity but also local functional groups and 3D geometry with high fidelity.
Figure 3: Visualization of molecular structure elucidation results using DiffSpectra under different configurations. We compared single-spectrum inputs (IR, Raman, UV-Vis), multi-modal spectra, and the effect of the pre-trained SpecFormer. Ground-truth structures are shown on the left for reference.
Table 2: Structure elucidation performance of DiffSpectra on the QM9S dataset. We compared two factors: (i) the effect of using a pre-trained versus an untrained SpecFormer as the spectral condition encoder, and (ii) the effect of generating molecular structures in 2D topology space versus in joint 2D&3D geometry space. Reported metrics include exact Acc@$K$, MCES, fingerprint- and fragment-based similarities, functional group similarity, and, where applicable, 3D-specific measures such as RMSD (in Å) and atom mapping accuracy.
Molecular Modality Pre-trained SpecFormer Acc@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$ $\uparrow$ $\operatorname{CosSim}_{\mathrm{MG}}$ $\uparrow$ $\operatorname{TaniSim}_{\mathrm{MA}}$ $\uparrow$ FraggleSim $\uparrow$ FGSim $\uparrow$ RMSD (Å) $\downarrow$ MapAcc $\uparrow$ 2D 7.95% 2.6581 0.5295 0.6478 0.8144 0.8999 0.8930 - - 9.49% 2.4132 0.5609 0.6734 0.8350 0.9072 0.9034 - - 2D&3D 35.64% 1.7795 0.7242 0.7964 0.8941 0.9397 0.9495 1.653 98.08% 40.76% 1.3337 0.7875 0.8450 0.9262 0.9504 0.9652 1.651 98.25%
@@ 2.4 Pre-trained SpecFormer facilitates more accurate structure elucidation
To assess the impact of pre-training the spectrum encoder, we conducted an ablation study comparing DiffSpectra with and without a pre-trained SpecFormer module, as shown in Figure 2, Figure 3 and Table 2. When equipped with the pre-trained SpecFormer, DiffSpectra achieves the top-1 accuracy of 40.76%, compared to 35.64% without pre-training, indicating a clear improvement in correctly recovering the ground-truth molecular structures from spectra. Similarly, the MCES decreases from 1.7795 to 1.3337, suggesting that molecular graphs predicted with pre-trained SpecFormer retain more of the correct connectivity of the target structures.
In addition, the fingerprint-based similarity scores also improve with pre-training. For example, the Tanimoto similarity over Morgan fingerprints rises from 0.7242 to 0.7875, and the cosine similarity improves from 0.7964 to 0.8450. The Tanimoto similarity over MACCS keys also increases from 0.8941 to 0.9262. These gains indicate that the molecular structures generated with pre-trained SpecFormer more accurately capture both local substructural patterns and functional-group motifs. Moreover, the Fraggle similarity improves from 0.9397 to 0.9504, and the FGSim from 0.9495 to 0.9652, confirming that larger chemical fragments and functional groups are better preserved.
@@ 2.5 Multi-modal spectra outperform single-modality spectra for structure elucidation
Among individual modalities, Raman spectra alone outperform IR and UV-Vis, with the top-1 accuracy of 30.3% and reasonable similarity scores. IR spectra alone still provide useful chemical information, achieving 25.9% top-1 accuracy, while UV-Vis alone performs poorly (0.3% top-1 accuracy), reflecting its limited ability to uniquely identify molecular structures in the QM9S dataset.
Overall, these results highlight that combining multiple spectral modalities provides complementary structural priors, enabling DiffSpectra to generate molecular structures with higher fidelity and greater consistency with ground-truth targets.
@@ 2.6 Sampling multiple candidates improves structural hit accuracy
Owing to the inherently stochastic nature of diffusion models, each sampling process may yield a distinct yet plausible molecular structure, even under the same conditions. Although a single sample may not always match the ground-truth structure exactly, generating multiple candidates increases the likelihood that at least one of them will align with the correct structure. This motivates the use of top-$K$ Accuracy (Acc@$K$) as a more comprehensive evaluation metric, formally defined in Section C.2.
@@
We report in Figure 2(C) the performance of different model variants across varying sample sizes. Specifically, Acc@$K$ quantifies the fraction of test molecules for which the correct structure appears within the top $K$ generated candidates. For instance, the top-1 accuracy corresponds to the exact-match rate under single-sample generation, whereas Acc@$K$ ($K>1$) reflects the probability of recovering the correct structure when multiple guesses are permitted. Across all model variants, Acc@$K$ exhibits a consistent monotonic increase with $K$, as the likelihood of including the correct structure naturally improves when more outputs are considered. Notably, the full model, when equipped with all spectra and pre-trained SpecFormer initialization, improves from 40.76% at $K=1$ to 99.49% at $K=10$. This demonstrates that even modest sampling budgets can yield near-exhaustive coverage of the true structure space.
These findings underscore an important property of diffusion-based molecular elucidation: although the model may not always recover the exact structure in a single attempt, it reliably ranks the correct structure within a compact set of plausible candidates. This property is particularly advantageous in practical workflows, where a ranked shortlist of candidates can be provided for subsequent experimental validation.
2.7 Generating 3D geometries helps achieve more precise elucidation
To assess the advantage of generating molecules in 3D geometry space, we compared structure elucidation in a purely 2D modality, where molecules are generated as topological graphs without spatial coordinates, against joint 2D and 3D generation. The results are summarized in Table 2.
Incorporating 3D geometries yields substantial gains. This improvement arises because molecular spectra are intrinsically more correlated with 3D structures than with 2D topology: spectra reflect energy absorption patterns, whereas 3D structures more faithfully capture the energetic and spatial states of molecules. Consequently, the joint 2D and 3D generation achieves markedly higher exact elucidation accuracy (the top-1 accuracy increases from 9.49% to 40.76%), reduced MCES, and consistent improvements across fingerprint, fragment, and functional group similarities. By contrast, RMSD and mapping accuracy cannot be computed under the 2D-only settings, as 3D coordinates are absent.
Moreover, even within the 2D modality, the use of a pre-trained SpecFormer as the spectral condition encoder leads to measurable gains (e.g., the top-1 accuracy improves from 7.95% to 9.49%), highlighting the benefit of spectral pre-training.
@@ 2.7 Generating 3D geometries helps achieve more precise elucidation
Table 3: Evaluation of different SE(3) equivariance strategies for DiffSpectra. We compared a model-based equivariant architecture to data-based equivariance approaches, with or without data augmentation, on molecular structure elucidation metrics. Here, “data aug” refers to data augmentation.
@@
SE(3) Equivariance Acc@1 $\uparrow$ MCES $\downarrow$ $\operatorname{TaniSim}_{\mathrm{MG}}$ $\uparrow$ $\operatorname{CosSim}_{\mathrm{MG}}$ $\uparrow$ $\operatorname{TaniSim}_{\mathrm{MA}}$ $\uparrow$ FraggleSim $\uparrow$ FGSim $\uparrow$ model-based 35.64% 1.7795 0.7242 0.7964 0.8941 0.9397 0.9495 data-based (w/ data aug) 31.24% 1.9571 0.7025 0.7803 0.8834 0.9359 0.9420 data-based (w/o data aug) 20.25% 3.6036 0.5138 0.6308 0.7588 0.8658 0.8620
@@ 2.8 Comparing model-based and data-based SE(3) equivariance strategies
Since our model performs diffusion generation in molecular 3D space, it is essential to ensure SE(3) equivariance so that predictions remain consistent under rigid-body transformations of the input. To this end, we designed and evaluated two strategies for achieving SE(3) equivariance: model-based equivariance and data-based equivariance. In the model-based approach, geometric inductive biases are explicitly incorporated into the network architecture, for example, through pairwise distance encodings and equivariant coordinate updates, which ensure equivariant behavior. By contrast, the data-based approach does not include SE(3)-equivariant inductive biases in the model architecture. Instead, it relies on random rotation and translation augmentations during training, encouraging the model to learn equivariance empirically from data.
@@
As shown in Table 3, the model-based SE(3)-equivariant implementation of DiffSpectra achieves consistently superior performance. It attains the highest top-1 accuracy (35.64%) and the lowest MCES (1.7795), indicating stronger recovery of molecular graph connectivity. The data-based approach with data augmentation achieves slightly lower top-1 accuracy (31.24%). Notably, when data augmentation is removed from the data-based approach, performance degrades substantially across all metrics, confirming the critical role of data augmentation in enabling the model to generalize over molecular rotations and translations. Overall, these results demonstrate that while data-based equivariance with augmentation can help achieve some degree of SE(3)-aware behavior, explicitly incorporating geometric inductive biases through model-based equivariance constitutes a more robust and effective strategy for molecular structure elucidation from spectra.
2.9 Generalization across molecules with different numbers of atoms
Figure 4: Molecular structure elucidation performance across molecules with varying numbers of atoms. The figure presents results obtained using different similarity metrics, including $\operatorname{TaniSim}_{\mathrm{MG}}$, $\operatorname{CosSim}_{\mathrm{MG}}$, $\operatorname{TaniSim}_{\mathrm{MA}}$, FraggleSim, and FGSim.
To evaluate the generalizability of DiffSpectra across molecules of varying sizes, we analyzed its performance on molecules containing different numbers of atoms. Figure 4 presents the results across multiple similarity metrics, where dashed lines denote raw similarity values and solid curves represent the smoothed trends obtained using the Savitzky-Golay filter [46].
As shown in Figure 4, the similarity metrics based on molecular fingerprints, such as $\mathrm{TaniSim}_{\mathrm{MG}}$, $\mathrm{CosSim}_{\mathrm{MG}}$, and $\mathrm{TaniSim}_{\mathrm{MA}}$, remain consistently stable across molecules of different sizes. This indicates that DiffSpectra achieves robust generalization regardless of molecular size. In contrast, fragment- and functional group-based similarity metrics (FraggleSim and FGSim) exhibit slightly lower performance when the number of atoms is small. However, as the number of atoms increases, both metrics steadily improve and approach values close to 1.0. This trend arises because very small molecules cannot be effectively decomposed into representative fragments or functional groups, which constrains the reliability of these similarity measures. With larger molecular structures, the decomposition becomes more informative, resulting in improved and stable performance.
In summary, these results demonstrate that DiffSpectra maintains reliable generalization across molecular scales, ensuring consistent performance for molecules of different sizes.
2.10 Elucidation trajectory analysis based on gradient
Figure 5: Gradient-based elucidation trajectory analysis. Gradient heatmaps depict atomic gradients with respect to spectral patches, with each heatmap averaged over a time interval of 0.05 (corresponding to 50 denoising steps). Upon completion of the denoising process, a molecular structure is elucidated that matches the ground-truth structure. Atom indices are annotated on the structure.
To further reveal the generative trajectory of DiffSpectra and gain deeper insight into its structural elucidation process, we analyzed the gradients of each atom with respect to the input spectral patches. Specifically, we computed the gradient of the predicted atom type for each atom with respect to the input patches at each denoising step. Since the denoising process in DiffSpectra consists of 1,000 steps, this procedure yields a total of 1,000 gradient heatmaps. In these heatmaps, each row corresponds to an indexed atom, while each column corresponds to a spectral patch. Larger gradient values indicate that, at the corresponding time step, a spectral patch exerts a stronger influence on the prediction of the atom’s type. To better visualize the generative trajectory, we averaged the gradients over every 50 steps, corresponding to time intervals of 0.05 along the generative time axis (ranging from 1 to 0). The resulting averaged gradient heatmaps, together with the final molecular structure annotated by atom indices, are presented in Figure 5.
Inspection of these heatmaps reveals that the generative behavior of DiffSpectra evolves dynamically throughout the denoising process. Based on these observations, we delineate the trajectory into four characteristic phases: initial, early, late, and final phases. During the initial phase (approximately from 1.0 to 0.7), the gradients of all atoms are elevated, suggesting that the model is establishing a broad semantic representation that provides the foundation for subsequent, more fine-grained assignments. During the early phase (approximately from 0.7 to 0.3), atoms within simple substructures, particularly  – C – C –  and  – C – H bonds, exhibit stronger gradients, indicating that these elementary motifs are elucidated first. The late phase (approximately from 0.3 to 0.1) is marked by the preferential elucidation of more complex functional groups, such as ethers ( – C – O – C – ), hydroxyls ( – OH), and alkynes ( – C –  –  – C – ), which are assigned higher gradient values during this phase. Finally, in the final phase (approximately from 0.1 to 0.0), the gradients associated with all atoms diminish gradually, signifying that DiffSpectra has largely completed the elucidation process and converged on a confident molecular structure.
In summary, gradient-based trajectory analysis reveals that DiffSpectra first generates initial semantics encompassing all atoms, then elucidates simple substructures, subsequently focuses on more complex functional groups, and finally consolidates the molecular structure with high confidence.